STEP 1 Force: Just after releasing the block, calculate the force exerted on the block by the spring, the force exerted on the block by the Earth, and the net force: spring,r# 2.64 Earthy-6769 net.y - 1.96 0.676N 1.96 N Momentum update: Just after releasing the block, the momentum of the block is zero. Approximate the average net force during the next timel by the force you just calculated. At t0.07 seconds, what will the new momentum and velocity of the block be? Py-1372 Vy1.98 kg m/s m/s Position update: Initially the bottom of the block is at y 0.12 m. Approximating the average velocity in the first time interval by the final velocity, what will be the new position of the bottom of the block at time t-0.07 seconds? r= .2592 STEP 2 Force: At the new position, calculate the force exerted on the block by the spring, the force exerted on the block by the Earth, and the net force spring.y FEarth,y Momentum update: Approximate the average net force during the next time interval by the force you just calculated. At time t = 2 x 0.07 = 0.14 seconds, what will the new momentum and velocity of the block be? kg m/s mys Position update: Approximating the average velocity in the second time interval by the final velocity, what will be the new position of the bottom of the block at time t-2 × 0.07-0.14 seconds? STEP 3 Force: At the new position, calculate the force exerted onthe block by the spring, the force exerted on the block by the Earth, and the net force: spring.y Earth.y Momentum update: Approximate the average net force during the next time interval by the force you just calculated. At time t-3 x 0.07-0.21 seconds, what will the new momentum and velocity of the block be? Py kg m/s mys Position update: Approximating the average velocity in the third time interval by the final velocity what will be the new position of the bottom of the block at time t = 3 x 0.07 0.21 seconds?Push down, Relaxed length release from rest A spring has a relaxed length of 34 cm (0.34 m) and its spring stiffness is 12 N/m. You glue a 69 gram block (0.069 kg) to the top of the spring, and push the block down, compressing the spring so its total length is 12 cm. You make sure the block is at rest, then at time t-0 you quickly move your hand away. The block begins to move upward, because the upward force on the block by the spring is greater than the downward force on the block by the Earth. Calculate approximately y vs. time for the block during a 0.21-second interval after you release the block, by applying the Momentum Principle in three steps each of 0.07-second duration. To avoid buildup of small errors causing you to lose credit, in Step 2 we use your answers to Step 1 even if they are not correct, and in Step 3 we use your answers to Step 2 even if they are not correct. We will only consider the y components in the following calculations, because there is no change in x or z
Answer #1

Step 2:

force exerted by the spring =spring constant*compression

=12*(0.34 - 0.2592) = 0.9696 N

force is directed vertically upwards.

part 2:

force by earth = weight of the block = mass*g

= 0.069*9.8 = 0.676 N

force is directed vertically downward.

part 3:

net force is in vertically upward direction.

net force = 0.9696 - 0.676 = 0.293 N

part 4:

change in momentum = force*time duration

py - 0.1372 = 0.293*0.07

py = 0.1577 kg.m/s

part 5:

velocity at t = 0.07 s

v = py/mass =0.1577/0.069 = 2.28 m/s

part 6:

new position at t=0.07 seconds

y = initial position+ velocity*time

y = 0.2592 + 2.28*0.07 = 0.4192 m

Step 3:

force exerted by the spring =spring constant*compression

=12*(0.34 - 0.4192) = -0.9504 N

force is directed vertically downward.

part 2:

force by earth = weight of the block = mass*g

=0.069*9.8 = 0.676 N

force is directed vertically downward.

part 3:

net force is in vertically upward direction.

net force = -0.9504 -0.676 = -1.626 N

part 4:

change in momentum = force*time duration

py - 0.1577 = -1.626*0.07

py = 0.0438 kg.m/s

part 5:

velocity at t=0.07 seconds

v = py/mass =0.0438/0.069 = 0.6347 m/s

part 6:

new position at t=0.04 seconds

y = initial position+ velocity*time

y =0.4192 +0.6347*0.07 = 0.4636 m